OpenStudy (anonymous):

If g(x) = sqrt(x-3) and k(x) = x^2 + 5 then determine: (g*k)(7)

OpenStudy (anonymous):

just multiply the two functions and evaluate at x=7

OpenStudy (anonymous):

How do you multiply them?

OpenStudy (anonymous):

(x-3) * (x^4 + 25). Is this what you are suppose to get when you multiply them?\[x ^{5}-3x ^{4}+25x-75\]

jimthompson5910 (jim_thompson5910):

you forgot the square root over the x-3

jimthompson5910 (jim_thompson5910):

(g*k)(7) is the same as g(7) * k(7)

OpenStudy (anonymous):

I squared the both equations so I can get rid of the square root.

jimthompson5910 (jim_thompson5910):

find g(7) and k(7). Then multiply those results

jimthompson5910 (jim_thompson5910):

squaring x^2 + 5 does not give x^4 + 25 also, squaring can only be done when you square both sides of an equation

OpenStudy (anonymous):

Would the answer be 216 then?

OpenStudy (anonymous):

Oh I keep forgetting some steps. But I got the answer doing what you said.

jimthompson5910 (jim_thompson5910):

it's not 216

OpenStudy (anonymous):

x^2 + 5 = 7^2 + 5 = 49 + 5 = 54 sqrt(x - 3) = sqrt(7 - 3) = sqrt(4) = 2 54 * 2 = 108 I forgot to sqrt the 4 earlier.

jimthompson5910 (jim_thompson5910):

108 is correct

OpenStudy (anonymous):

Thanks!

jimthompson5910 (jim_thompson5910):

no problem